Let u=(-5, -2, 4), v= (3, 6, -5) and w = (-7, 1, -8)

a) Calculate (u × v).w and hence find the volume of the parallelepiped with adjacent sides u, v, and w.

b) Show that vector (u - projwu) and w are orthogonal.

c) Use the cross product to find the angle between u and w.

1
Expert's answer
2022-05-24T23:22:14-0400

a)


"(\vec {u}\times\vec {v})\vec {w}=\begin{vmatrix}
 -5 & -2 & 4 \\
 3 & 6 & -5 \\
 -7 & 1 & -8 \\
\end{vmatrix}"

"=-5\begin{vmatrix}
 6 & -5 \\
 1 & -8
\end{vmatrix}+2\begin{vmatrix}
 3 & -5 \\
 -7 & -8
\end{vmatrix}+4\begin{vmatrix}
 3 & 6 \\
 -7 & 1
\end{vmatrix}"

"=-5(-48+5)+2(-24-35)+4(3+42)"

"=277"

"V=277" cubic units


b)


"proj_{\vec{w}}\vec{u}=\dfrac{\vec{u}\cdot\vec{w}}{|\vec{w}|^2}\vec{w}"

"\vec {u}-proj_{\vec{w}}\vec{u}=\vec {u}-\dfrac{\vec{u}\cdot\vec{w}}{|\vec{w}|^2}\vec{w}"

"(\vec {u}-proj_{\vec{w}}\vec{u})\cdot\vec {w}=(\vec {u}-\dfrac{\vec{u}\cdot\vec{w}}{|\vec{w}|^2}\vec{w})\cdot\vec {w}"

"=\vec{u}\cdot\vec{w}-\dfrac{\vec{u}\cdot\vec{w}}{|\vec{w}|^2}|\vec{w}|^2"




"=\vec{u}\cdot\vec{w}-\vec{u}\cdot\vec{w}=0"

Therefore "(\vec {u}-proj_{\vec{w}}\vec{u})" and "\vec{w}" are orthogonal.


c)



"\vec {u}\times\vec {w}=\begin{vmatrix}
 \vec {i} & \vec {j} & \vec {k} \\
 -5 & -2 & 4 \\
 -7 & 1 & -8 \\
\end{vmatrix}"

"=\vec {i}\begin{vmatrix}
 -2 & 4 \\
 1 & -8
\end{vmatrix}-\vec {j}\begin{vmatrix}
 -5 & 4 \\
 -7 & -8
\end{vmatrix}+\vec {k}\begin{vmatrix}
 -5 & -2 \\
 -7 & 1
\end{vmatrix}"

"=12\vec {i}-68\vec {j}-19\vec {k}"


"|\vec{u}\times\vec{w}|=\sqrt{(12)^2+(-68)^2+(-19)^2}"

"=\sqrt{5129}"

"|\vec{u}|=\sqrt{(-5)^2+(-2)^2+(4)^2}=\sqrt{47}"

"|\vec{w}|=\sqrt{(-7)^2+(1)^2+(-8)^2}=\sqrt{114}"

"\sin \theta=\dfrac{|\vec{u}\times\vec{w}|}{|\vec{u}||\vec{w}|}=\sqrt{\dfrac{5129}{47(114)}}\approx0.9784"

"\theta=\sin^{-1}{\sqrt{\dfrac{5129}{47(114)}}}\approx78\degree"